OpenStudy (anonymous):

One positive integer is 2 less than another. The product of the two integers is 24. What are the integers? Separate the answers with a comma.

OpenStudy (anonymous):

x=y-2 xy=24 (y-2)y=24 Now find y, then plug it into x=y-2 to find x.

OpenStudy (anonymous):

Howdoyoufind y?

OpenStudy (anonymous):

@alexsmith4589

OpenStudy (anonymous):

oh heck, i would just make a guess

OpenStudy (anonymous):

what two numbers multiply to 24?

OpenStudy (anonymous):

12 and 2 do, but 2 is not 2 less than 12, skip that one

OpenStudy (anonymous):

how about 6 and 4?

OpenStudy (anonymous):

(y-2)y=24 y*y - y*(-2) = 24 y^2-2y=24 y^2-2y-24=0 solve y from that, I'm not sure how to show it

OpenStudy (anonymous):

yup, \(6\times 4=24\) and \(6-4=2\) got it on the second try

OpenStudy (anonymous):

I can't. I'm in online school, anytime I get anyyyyyything wrong my gpa goes down by a ridiculous amount

OpenStudy (anonymous):

i didn't mean guess the answer and just put it in i meant guess what it might be and check to see if it is right

OpenStudy (anonymous):

there are not too many choices for whole numbers whose product is 24

OpenStudy (anonymous):

Ahh, okay. I don't understand how alexander got y but the guessing then testing worked, thanks guys :)